Q3- Assuming monthly payments, which one of the following plans are better for buying a $15,000 van?
a) 6% on the full amount for 4 years.
b) $2500 discount and 12% interest on the remaining amount for 4 years.

please explain all steps and show equations. thank you
Answer #1

a) No of Months = 12* 4 = 48
Rate per month = 6%/12
Periodic Payments = 15000/(1-(1+0.5%)-48)/0.5% = 352.28

b)PV = 15000 - 2500 = 12500
Rate per month = 12/12% = 1%
Periodic Payments = 12500/(1-(1+1%)-48)/1% = 329.17

Second payment is better.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well